# Linear-scaling Tensor Train Sketching

### 11 Mar 2026 v1

@@ Abstract
We introduce the Block Sparse Tensor Train (BSTT) sketch, a structured random projection tailored to the tensor train (TT) format that unifies existing TT-adapted sketching operators. By varying two integer parameters $P$ and $R$, BSTT interpolates between the Khatri-Rao sketch ($R=1$) and the Gaussian TT sketch ($P=1$). We prove that BSTT satisfies an oblivious subspace embedding (OSE) property with parameters $R=\mathcal{O}(d(r+\log 1/\delta))$ and $P=\mathcal{O}(\varepsilon^{-2})$, and an oblivious subspace injection (OSI) property under the condition $R=\mathcal{O}(d)$ and $P=\mathcal{O}(\varepsilon^{-2}(r+\log r/\delta))$. Both guarantees depend only linearly on the tensor order $d$ and on the subspace dimension $r$, in contrast to prior constructions that suffer from exponential scaling in $d$. As direct consequences, we derive quasi-optimal error bounds for the QB factorization and randomized TT rounding. The theoretical results are supported by numerical experiments on synthetic tensors, Hadamard products, and a quantum chemistry application.
@@ 1 Introduction
In this work, we introduce a family of tensor train sketches, unifying the previously known classes into one that we call the Block-Sparse Tensor Train (BSTT) sketch, and for which we provide new theoretical guarantees such as an oblivious subspace embedding with parameters that depend polynomially on the dimension or the ranks, and an oblivious subspace injection with less stringent conditions on the parameters than its embedding counterpart. Furthermore, we apply these results to show that the randomized TT rounding algorithm and low-rank approximation output a quasi-optimal tensor train.
The remainder of this work is organized as follows. Section 2 outlines the notation used throughout the paper and presents the probabilistic setting used in our main results. Section 3 presents our core contributions, introducing the new BSTT sketching framework and its theoretical guarantees, alongside its practical implications for low-rank approximation and randomized tensor train rounding algorithms. Section 4 provides a comparative self-contained review of existing sketching operators for the tensor train format, highlighting their respective sample complexities and inherent limitations. Section 5 details the performed numerical experiments where the BSTT sketch is applied to different scenarios that require randomized rounding, thus providing empirical evidence that small parameters are sufficient for practical applications. Section 6 outlines the structural properties necessary to prove that BSTT acts as an oblivious subspace embedding. Subsequently, Section 7 develops the novel analytical machinery required to establish that BSTT satisfies the oblivious subspace injection property, and concludes with the formal error analysis for randomized TT rounding and low-rank approximation.
@@ 3 Main Results
In this section, we present the theoretical contributions of our work, establishing probabilistic guarantees for the proposed BSTT sketching framework and its use in applications. Proofs of the embedding results are postponed to Sections 7 and 6.
3.1 Block-Sparse Tensor Train Sketch
@@
Definition 3.1 (Block-Sparse Tensor Train Sketch).
Given integer parameters $P,R\in\mathbb{N}$, the Block-Sparse Tensor Train sketch matrix ${\bm{\Omega}}_{\mathtt{BSTT}}\in\mathbb{F}^{PR\times N}$ with tensor-structured row blocks is defined as
${\bm{\Omega}}_{\mathtt{BSTT}}:=\frac{1}{\sqrt{P}}\begin{bmatrix}\bigl(\mathcal{G}^{(1,1)}\bowtie\cdots\bowtie\mathcal{G}^{(1,d)}\bigr)^{\leq 1}\\ \vdots\\ \bigl(\mathcal{G}^{(P,1)}\bowtie\cdots\bowtie\mathcal{G}^{(P,d)}\bigr)^{\leq 1}\end{bmatrix}$, (4)
@@
We can readily observe that the BSTT sketch constitutes a parameterized family of sketching operators governed by the inputs $P$ and $R$. Notably, this formulation interpolates between other well-known tensor-adapted sketches: the Khatri-Rao sketch [43, 12] when $R=1$, and the random Gaussian TT-tensor sketch [4, 26, 22] when $P=1$.
@@
We call this sketch ”Block-Sparse Tensor Train” because it may be reformulated as a contraction of a single block tensor train with a left boundary rank of dimension $PR$, by writing ${\bm{\Omega}}_{\mathtt{BSTT}}=\frac{1}{\sqrt{P}}(\mathcal{G}_{1}\bowtie\cdots\bowtie\mathcal{G}_{d})^{\leq 1}$, where the mode-wise slices of each core have the block-diagonal structure:
@@
We note that the BSTT sketch is a structured sketch of embedding dimension $PR$ whose application cost is equivalent to contracting $P$ tensor trains of rank $R$ with the input tensor train of maximum TT-rank $\chi$. Thus, to sketch a tensor train of rank $\chi$ with BSTT, the cost scales as $\mathcal{O}(dnPR\chi(\chi+R))$, which is only modestly larger than with the Khatri-Rao sketch. Section 5.2 gives further details on this topic, emphasizing further optimizations in the case of structured input cores occurring in linear combination, Hadamard products and matrix-vector products for vectors and operators in TT format.
@@
Definition 3.4 (Orthogonal Block-Sparse Tensor Train Sketch).
Given integer parameters $P,R\in\mathbb{N}$, the Orthogonal Block-Sparse Tensor Train sketch matrix ${\bm{\Omega}}_{\mathtt{OBSTT}}\in\mathbb{F}^{PR\times N}$ with tensor-structured row blocks is defined as
${\bm{\Omega}}_{\mathtt{OBSTT}}:=\frac{1}{\sqrt{P}}\begin{bmatrix}\bigl(\sqrt{\frac{\rho_{1}n_{1}}{\rho_{0}}}\mathcal{U}^{(1,1)}\bowtie\cdots\bowtie\sqrt{\frac{\rho_{d}n_{d}}{\rho_{d-1}}}\mathcal{U}^{(1,d)}\bigr)^{\leq 1}\\ \vdots\\ \bigl(\sqrt{\frac{\rho_{1}n_{1}}{\rho_{0}}}\mathcal{U}^{(P,1)}\bowtie\cdots\bowtie\sqrt{\frac{\rho_{d}n_{d}}{\rho_{d-1}}}\mathcal{U}^{(P,d)}\bigr)^{\leq 1}\end{bmatrix}$, (5)
@@
Because of (3), ${\bm{\Omega}}_{\mathtt{OBSTT}}$ is effectively drawn uniformly from the Stiefel manifold in $\mathbb{F}^{PR\times N}$ with ${\bm{\Omega}}_{\mathtt{OBSTT}}{\bm{\Omega}}_{\mathtt{OBSTT}}^{*}=\frac{N}{PR}\mathbf{I}$. In addition, constructing ${\bm{\Omega}}_{\mathtt{OBSTT}}$ amounts to choosing each tensor train on the block-structured rows as a right-orthogonalized block tensor train. The improved injectivity and dilation ratios compared to the plain ${\bm{\Omega}}_{\mathtt{BSTT}}$ sketch in numerical experiments may be a consequence of this fact, but a rigorous study is left to future work.
@@
BSTT (This work) $m=PR$ $R=O\big(d(r+\log(1/\delta))\big)$ $P=\mathcal{O}(\varepsilon^{-2})$ $R=\mathcal{O}(d)$ $P=O\big(\varepsilon^{-2}(r+\log(r/\delta))\big)$ $\mathcal{O}\bigl(dnPR\chi(R+\chi)\bigr)$
Table 1: Comparison of probabilistic guarantees for tensor-train-adapted sketches. $n$ denotes the maximum mode dimension, $P$ denotes the number of independent entries or blocks in the Khatri-Rao, $f_{\mathtt{TT}(R)}$ or BSTT sketches; $R$ the TT-rank of random $T$ cores; and $\chi$ the maximum TT-rank of tensor trains to be sketched.
@@ 3.2.1 OSE guarantees
We begin by establishing sufficient conditions on the parameters $P$ and $R$ for the BSTT sketch to satisfy the OSE property. We do not pursue such bounds for the OBSTT sketch in this work, but numerical experiments indicate it outperforms the plain BSTT sketch. Section 4 places these bounds in the context of existing constructions.
@@
The Block-Sparse Tensor Train sketch is an $(\varepsilon,\delta,r)$-OSE provided
@@
Figure 1: BSTT Framework on Rank-1 Basis. Empirical injectivity (left) and dilation (right) of the proposed BSTT sketches. The operators are evaluated using embedding dimensions $PR=2r$ and block ranks $R\in\{1,4,16,32\}$, applied to $r$-dimensional target subspaces spanned by Kronecker (rank-1) Gaussian TT vectors in $(\mathbb{R}^{4})^{\otimes d}$. Markers indicate the median across 100 independent trials, with shaded regions denoting the interquartile range (25th to 75th percentiles).
Figure 2: BSTT Framework on Rank-4 Basis. Empirical injectivity (left) and dilation (right) of the proposed BSTT sketches. The operators are evaluated using embedding dimensions $PR=2r$ and block ranks $R\in\{1,4,16,32\}$, applied to $r$-dimensional target subspaces spanned by rank-44 Gaussian TT vectors in $\mathbb{R}^{4^{d}}$. Markers indicate the median across 100 independent trials, with shaded regions denoting the interquartile range (25th to 75th percentiles).
Experiments reported in Figures 1 and 2 evaluate the OSE property of the BSTT sketching framework. Specifically, we estimate the injectivity and dilation constants as functions of $d$ and $r$, for target subspaces spanned by rank-1 and rank-4 Gaussian TT vectors. While all block-structured sketches (particularly for $R\geq 16$) achieve a controlled embedding quality, this control tightens significantly as the TT-ranks of the basis increase. Notably, however, the exponential degradation of Khatri-Rao embeddings dissipates as the structure of the target subspace becomes more entangled – a concept explored in more detail in Section 7.
Finally, we establish probabilistic guarantees for BSTT to enable the randomized singular value decomposition (or randomized rangefinder).
@@
Let $\mathbf{A}\in\mathbb{F}^{k\times N}$ be an input matrix, $r\leq k$ the target rank, $\mathbf{A}_{r}$ the best rank-$r$ approximation of $\mathbf{A}$. There exists a constant $c>0$ such that given a Block-Sparse Tensor Train matrix ${\bm{\Omega}}_{\mathtt{BSTT}}\in\mathbb{F}^{PR\times N}$ satisfying for some $(\varepsilon,\delta)\in(0,1)$
@@ 3.2.2 OSI guarantees
While the OSE property provides the strongest geometric guarantee, it often imposes conservative bounds on the sketch parameters. With the analysis developed in Section 7, we show that BSTT satisfies the OSI property under milder conditions on the block count $P$ and rank $R$, and we further identify a notion of subspace entanglement $C_{\mathbf{Q}}(R)$ (Definition 7.9). The constant $C_{\mathbf{Q}}(R)$ captures the dependence of the embedding error on Kronecker-structured (or low-entanglement) vectors in the subspace spanned by $\mathbf{Q}$. Such vectors are precisely those for which the Khatri-Rao sketch degrades exponentially, a phenomenon termed overwhelming orthogonality [8, 30].
@@
Fix an orthonormal matrix $\mathbf{Q}\in\mathbb{F}^{N\times r}$. There exists a measure of subspace entanglement $0<C_{\mathbf{Q}}(R)\leq\left(1+\frac{p_{\mathbb{F}}}{R}\right)^{d}-1$, with $p_{\mathbb{R}}=2$ or $p_{\mathbb{C}}=1$ depending on the choice of scalar field $\mathbb{F}$, such that a random Block-Sparse Tensor Train matrix ${\bm{\Omega}}_{\mathtt{BSTT}}\in\mathbb{F}^{PR\times N}$ satisfies $\sigma_{min}^{2}({\bm{\Omega}}_{\mathtt{BSTT}}\mathbf{Q})\geq 1-\varepsilon$ with probability at least $1-\delta$ provided
@@
In particular, the BSTT matrix enjoys the $(1-\varepsilon,\delta,r)$-OSI property provided $R\sim p_{\mathbb{F}}d$ and $P=\mathcal{O}(\varepsilon^{-2}(r+\log(r/\delta)))$.
@@
Table 1 summarizes the comparative analysis described in Section 4, showing that while the Khatri-Rao sketch necessarily exhibits an exponential scaling in $d$ to achieve an OSE/OSI [2, 8], the Block-Sparse Tensor Train sketch achieves an OSE for block ranks $R=\mathcal{O}(d\log(1/\delta)$, or an OSI for $R=\mathcal{O}(d)$, whenever the number of blocks $P$ satisfies the conditions of Theorems 3.7 and 3.10 respectively. Numerical experiments conducted on the empirical injectivity and dilation $\sigma_{min/max}^{2}({\bm{\Omega}}_{\mathtt{(O)BSTT}}\mathbf{Q})$ for BSTT embeddings with block rank $R=2d$ applied to subspaces of dimension 16 for $d\leq 100$, presented in Figure 3, confirm this phenomenon. In particular, the injectivity parameter seems asymptotically constant, $\alpha=\mathcal{O}(1)$, and exhibits relatively small variance. The distortion parameter on the other hand exhibits larger variance (especially for $P=1$). Theory indicates its distribution is quite heavy-tailed, as it behaves conceptually as a product of $d$ chi-squared distributed variables. In all cases, the Orthogonal BSTT variant has better performance than the basic BSTT sketch, and increasing $P$ clearly reduces the embedding distortion.
Figure 3: Numerical OSE of BSTT Framework on Rank-1 Basis Empirical injectivity (left) and dilation (right) of the proposed BSTT sketches. The operators are evaluated using embedding dimensions PRPR with $P\in\{1,4,16\}$ blocks of rank $R=2d$, applied to a $16$-dimensional target subspace spanned by Kronecker (rank-1) Gaussian TT vectors in $(\mathbb{R}^{4})^{\otimes d}$. Markers indicate the median across 100 independent trials, with shaded regions denoting the interquartile range (25th to 75th percentiles).
As a direct consequence of Theorem 3.10, we establish probabilistic guarantees for the randomized rangefinder implemented with the BSTT sketch in the OSI parameter regime.
@@
Let $\mathbf{A}\in\mathbb{F}^{k\times N}$ be an input matrix, $r\leq k$ the target rank, $\mathbf{A}_{r}$ the best rank-$r$ approximation of $\mathbf{A}$, and ${\bm{\Omega}}_{\mathtt{BSTT}}\in\mathbb{F}^{PR\times N}$ a Block-Sparse Tensor Train sketch realizing an $(\alpha,\delta/2,r)$-OSI of the subspace spanned by the dominant $r$ singular vectors of $\mathbf{A}$ as prescribed by (7). Then, the rank-$PR$ QB factorization of $\widehat{\mathbf{A}}\in\mathbb{F}^{k\times N}$ determined by
@@
If we require that the sketch ${\bm{\Omega}}_{\mathtt{BSTT}}$ satisfies a Strong $(\alpha\sqrt{\varepsilon/r},\delta/2)$-Johnson-Lindenstrauss Moment property, Definition 6.1, then we may achieve an $(1+\varepsilon)$-accurate randomized SVD with probability $1-\delta$, thanks to Imported Lemma 6.10. This increases the sketching rank conditions to
@@
Finally, we address the application of our framework to operations intrinsic to the tensor train format. An important implication of the property established by Theorem 3.10 is the quality of the BSTT embedding for its application to the Randomized TT-Rounding. The following theorem demonstrates that our sketch produces a rounding approximation that preserves the quasi-optimality of the randomized SVD.

### 13 Mar 2026 v2

@@ Abstract
We introduce the TTStack sketch, a structured random projection tailored to the tensor train (TT) format that unifies existing TT-adapted sketching operators. By varying two integer parameters $P$ and $R$, TTStack interpolates between the Khatri-Rao sketch ($R=1$) and the Gaussian TT sketch ($P=1$). We prove that TTStack satisfies an oblivious subspace embedding (OSE) property with parameters $R=\mathcal{O}(d(r+\log 1/\delta))$ and $P=\mathcal{O}(\varepsilon^{-2})$, and an oblivious subspace injection (OSI) property under the condition $R=\mathcal{O}(d)$ and $P=\mathcal{O}(\varepsilon^{-2}(r+\log r/\delta))$. Both guarantees depend only linearly on the tensor order $d$ and on the subspace dimension $r$, in contrast to prior constructions that suffer from exponential scaling in $d$. As direct consequences, we derive quasi-optimal error bounds for the QB factorization and randomized TT rounding. The theoretical results are supported by numerical experiments on synthetic tensors, Hadamard products, and a quantum chemistry application.
@@ 1 Introduction
In this work, we introduce a family of tensor train sketches, unifying the previously known classes into one that we call the TTStack (TTS) sketch, and for which we provide new theoretical guarantees such as an oblivious subspace embedding with parameters that depend polynomially on the dimension or the ranks, and an oblivious subspace injection with less stringent conditions on the parameters than its embedding counterpart. Furthermore, we apply these results to show that the randomized TT rounding algorithm and low-rank approximation output a quasi-optimal tensor train.
The remainder of this work is organized as follows. Section 2 outlines the notation used throughout the paper and presents the probabilistic setting used in our main results. Section 3 presents our core contributions, introducing the new TTStack sketching framework and its theoretical guarantees, alongside its practical implications for low-rank approximation and randomized tensor train rounding algorithms. Section 4 provides a comparative self-contained review of existing sketching operators for the tensor train format, highlighting their respective sample complexities and inherent limitations. Section 5 details the performed numerical experiments where the TTStack sketch is applied to different scenarios that require randomized rounding, thus providing empirical evidence that small parameters are sufficient for practical applications. Section 6 outlines the structural properties necessary to prove that TTStack acts as an oblivious subspace embedding. Subsequently, Section 7 develops the novel analytical machinery required to establish that TTStack satisfies the oblivious subspace injection property, and concludes with the formal error analysis for randomized TT rounding and low-rank approximation.
@@ 3 Main Results
In this section, we present the theoretical contributions of our work, establishing probabilistic guarantees for the proposed TTStack sketching framework and its use in applications. Proofs of the embedding results are postponed to Sections 7 and 6.
3.1 TTStack Sketch
@@
Definition 3.1 (TTStack Sketch).
Given integer parameters $P,R\in\mathbb{N}$, the TTStack sketch matrix ${\bm{\Omega}}_{\mathtt{TTS}}\in\mathbb{F}^{PR\times N}$ with tensor-structured row blocks is defined as
${\bm{\Omega}}_{\mathtt{TTS}}:=\frac{1}{\sqrt{P}}\begin{bmatrix}\bigl(\mathcal{G}^{(1,1)}\bowtie\cdots\bowtie\mathcal{G}^{(1,d)}\bigr)^{\leq 1}\\ \vdots\\ \bigl(\mathcal{G}^{(P,1)}\bowtie\cdots\bowtie\mathcal{G}^{(P,d)}\bigr)^{\leq 1}\end{bmatrix}$, (4)
@@
We can readily observe that the TTStack sketch constitutes a parameterized family of sketching operators governed by the inputs $P$ and $R$. Notably, this formulation interpolates between other well-known tensor-adapted sketches: the Khatri-Rao sketch [43, 12] when $R=1$, and the random Gaussian TT-tensor sketch [4, 26, 22] when $P=1$.
@@
We call this sketch ”TTStack” because it may be reformulated as a contraction of a single block tensor train with a left boundary rank of dimension $PR$, by writing ${\bm{\Omega}}_{\mathtt{BSTT}}=\frac{1}{\sqrt{P}}(\mathcal{G}_{1}\bowtie\cdots\bowtie\mathcal{G}_{d})^{\leq 1}$, where the mode-wise slices of each core have the block-diagonal structure:
@@
We note that the TTStack sketch is a structured sketch of embedding dimension $PR$ whose application cost is equivalent to contracting $P$ tensor trains of rank $R$ with the input tensor train of maximum TT-rank $\chi$. Thus, to sketch a tensor train of rank $\chi$ with TTStack, the cost scales as $\mathcal{O}(dnPR\chi(\chi+R))$, which is only modestly larger than with the Khatri-Rao sketch. Section 5.2 gives further details on this topic, emphasizing further optimizations in the case of structured input cores occurring in linear combination, Hadamard products and matrix-vector products for vectors and operators in TT format.
@@
Definition 3.4 (Orthogonal TTStack Sketch).
Given integer parameters $P,R\in\mathbb{N}$, the Orthogonal TTStack sketch matrix ${\bm{\Omega}}_{\mathtt{OTTS}}\in\mathbb{F}^{PR\times N}$ with tensor-structured row blocks is defined as
${\bm{\Omega}}_{\mathtt{OTTS}}:=\frac{1}{\sqrt{P}}\begin{bmatrix}\bigl(\sqrt{\frac{\rho_{1}n_{1}}{\rho_{0}}}\mathcal{U}^{(1,1)}\bowtie\cdots\bowtie\sqrt{\frac{\rho_{d}n_{d}}{\rho_{d-1}}}\mathcal{U}^{(1,d)}\bigr)^{\leq 1}\\ \vdots\\ \bigl(\sqrt{\frac{\rho_{1}n_{1}}{\rho_{0}}}\mathcal{U}^{(P,1)}\bowtie\cdots\bowtie\sqrt{\frac{\rho_{d}n_{d}}{\rho_{d-1}}}\mathcal{U}^{(P,d)}\bigr)^{\leq 1}\end{bmatrix}$, (5)
@@
Because of (3), ${\bm{\Omega}}_{\mathtt{OTTS}}$ is effectively drawn uniformly from the Stiefel manifold in $\mathbb{F}^{PR\times N}$ with ${\bm{\Omega}}_{\mathtt{OTTS}}{\bm{\Omega}}_{\mathtt{OTTS}}^{*}=\frac{N}{PR}\mathbf{I}$. In addition, constructing ${\bm{\Omega}}_{\mathtt{OTTS}}$ amounts to choosing each tensor train on the block-structured rows as a right-orthogonalized block tensor train. The improved injectivity and dilation ratios compared to the plain ${\bm{\Omega}}_{\mathtt{TTS}}$ sketch in numerical experiments may be a consequence of this fact, but a rigorous study is left to future work.
@@
TTS (This work) $m=PR$ $R=O\big(d(r+\log(1/\delta))\big)$ $P=\mathcal{O}(\varepsilon^{-2})$ $R=\mathcal{O}(d)$ $P=O\big(\varepsilon^{-2}(r+\log(r/\delta))\big)$ $\mathcal{O}\bigl(dnPR\chi(R+\chi)\bigr)$
Table 1: Comparison of probabilistic guarantees for tensor-train-adapted sketches. $n$ denotes the maximum mode dimension, $P$ denotes the number of independent entries or blocks in the Khatri-Rao, $f_{\mathtt{TT}(R)}$ or TTStack sketches; $R$ the TT-rank of random $T$ cores; and $\chi$ the maximum TT-rank of tensor trains to be sketched.
@@ 3.2.1 OSE guarantees
We begin by establishing sufficient conditions on the parameters $P$ and $R$ for the TTStack sketch to satisfy the OSE property. We do not pursue such bounds for the Orthogonal TTStack sketch in this work, but numerical experiments indicate it outperforms the plain TTStack sketch. Section 4 places these bounds in the context of existing constructions.
@@
The TTStack sketch is an $(\varepsilon,\delta,r)$-OSE provided
@@
Figure 1: TTStack Framework on Rank-1 Basis. Empirical injectivity (left) and dilation (right) of the TTStack sketch. The operators are evaluated using embedding dimensions $PR=2r$ and block ranks $R\in\{1,4,16,32\}$, applied to $r$-dimensional target subspaces spanned by Kronecker (rank-1) Gaussian TT vectors in $(\mathbb{R}^{4})^{\otimes d}$. Markers indicate the median across 100 independent trials, with shaded regions denoting the interquartile range (25th to 75th percentiles).
Figure 2: TTStack Framework on Rank-4 Basis. Empirical injectivity (left) and dilation (right) of the TTStack sketches. The operators are evaluated using embedding dimensions $PR=2r$ and block ranks $R\in\{1,4,16,32\}$, applied to $r$-dimensional target subspaces spanned by rank-44 Gaussian TT vectors in $\mathbb{R}^{4^{d}}$. Markers indicate the median across 100 independent trials, with shaded regions denoting the interquartile range (25th to 75th percentiles).
Experiments reported in Figures 1 and 2 evaluate the OSE property of the TTStack sketching framework. Specifically, we estimate the injectivity and dilation constants as functions of $d$ and $r$, for target subspaces spanned by rank-1 and rank-4 Gaussian TT vectors. While all block-structured sketches (particularly for $R\geq 16$) achieve a controlled embedding quality, this control tightens significantly as the TT-ranks of the basis increase. Notably, however, the exponential degradation of Khatri-Rao embeddings dissipates as the structure of the target subspace becomes more entangled – a concept explored in more detail in Section 7.
Finally, we establish probabilistic guarantees for TTStack to enable the randomized singular value decomposition (or randomized rangefinder).
@@
Let $\mathbf{A}\in\mathbb{F}^{k\times N}$ be an input matrix, $r\leq k$ the target rank, $\mathbf{A}_{r}$ the best rank-$r$ approximation of $\mathbf{A}$. There exists a constant $c>0$ such that given a TTStack matrix ${\bm{\Omega}}_{\mathtt{TTS}}\in\mathbb{F}^{PR\times N}$ satisfying for some $(\varepsilon,\delta)\in(0,1)$
@@ 3.2.2 OSI guarantees
While the OSE property provides the strongest geometric guarantee, it often imposes conservative bounds on the sketch parameters. With the analysis developed in Section 7, we show that TTStack satisfies the OSI property under milder conditions on the block count $P$ and rank $R$, and we further identify a notion of subspace entanglement $C_{\mathbf{Q}}(R)$ (Definition 7.9). The constant $C_{\mathbf{Q}}(R)$ captures the dependence of the embedding error on Kronecker-structured (or low-entanglement) vectors in the subspace spanned by $\mathbf{Q}$. Such vectors are precisely those for which the Khatri-Rao sketch degrades exponentially, a phenomenon termed overwhelming orthogonality [8, 30].
@@
Fix an orthonormal matrix $\mathbf{Q}\in\mathbb{F}^{N\times r}$. There exists a measure of subspace entanglement $0<C_{\mathbf{Q}}(R)\leq\left(1+\frac{p_{\mathbb{F}}}{R}\right)^{d}-1$, with $p_{\mathbb{R}}=2$ or $p_{\mathbb{C}}=1$ depending on the choice of scalar field $\mathbb{F}$, such that a random TTStack matrix ${\bm{\Omega}}_{\mathtt{TTS}}\in\mathbb{F}^{PR\times N}$ satisfies $\sigma_{min}^{2}({\bm{\Omega}}_{\mathtt{TTS}}\mathbf{Q})\geq 1-\varepsilon$ with probability at least $1-\delta$ provided
@@
In particular, the TTStack matrix enjoys the $(1-\varepsilon,\delta,r)$-OSI property provided $R\sim p_{\mathbb{F}}d$ and $P=\mathcal{O}(\varepsilon^{-2}(r+\log(r/\delta)))$.
@@
Table 1 summarizes the comparative analysis described in Section 4, showing that while the Khatri-Rao sketch necessarily exhibits an exponential scaling in $d$ to achieve an OSE/OSI [2, 8], the TTStack sketch achieves an OSE for block ranks $R=\mathcal{O}(d\log(1/\delta)$, or an OSI for $R=\mathcal{O}(d)$, whenever the number of blocks $P$ satisfies the conditions of Theorems 3.7 and 3.10 respectively. Numerical experiments conducted on the empirical injectivity and dilation $\sigma_{min/max}^{2}({\bm{\Omega}}_{\mathtt{(O)TTS}}\mathbf{Q})$ for TTStack embeddings with block rank $R=2d$ applied to subspaces of dimension 16 for $d\leq 100$, presented in Figure 3, confirm this phenomenon. In particular, the injectivity parameter seems asymptotically constant, $\alpha=\mathcal{O}(1)$, and exhibits relatively small variance. The distortion parameter on the other hand exhibits larger variance (especially for $P=1$). Theory indicates its distribution is quite heavy-tailed, as it behaves conceptually as a product of $d$ chi-squared distributed variables. In all cases, the Orthogonal TTStack variant has better performance than the basic TTStack sketch, and increasing $P$ clearly reduces the embedding distortion.
Figure 3: Numerical OSE of TTStack Framework on Rank-1 Basis Empirical injectivity (left) and dilation (right) of the TTStack sketches. The operators are evaluated using embedding dimensions PRPR with $P\in\{1,4,16\}$ blocks of rank $R=2d$, applied to a $16$-dimensional target subspace spanned by Kronecker (rank-1) Gaussian TT vectors in $(\mathbb{R}^{4})^{\otimes d}$. Markers indicate the median across 100 independent trials, with shaded regions denoting the interquartile range (25th to 75th percentiles).
As a direct consequence of Theorem 3.10, we establish probabilistic guarantees for the randomized rangefinder implemented with the TTStack sketch in the OSI parameter regime.
@@
Let $\mathbf{A}\in\mathbb{F}^{k\times N}$ be an input matrix, $r\leq k$ the target rank, $\mathbf{A}_{r}$ the best rank-$r$ approximation of $\mathbf{A}$, and ${\bm{\Omega}}_{\mathtt{BSTT}}\in\mathbb{F}^{PR\times N}$ a TTStack sketch realizing an $(\alpha,\delta/2,r)$-OSI of the subspace spanned by the dominant $r$ singular vectors of $\mathbf{A}$ as prescribed by (7). Then, the rank-$PR$ QB factorization of $\widehat{\mathbf{A}}\in\mathbb{F}^{k\times N}$ determined by
@@
If we require that the sketch ${\bm{\Omega}}_{\mathtt{TTS}}$ satisfies a Strong $(\alpha\sqrt{\varepsilon/r},\delta/2)$-Johnson-Lindenstrauss Moment property, Definition 6.1, then we may achieve an $(1+\varepsilon)$-accurate randomized SVD with probability $1-\delta$, thanks to Imported Lemma 6.10. This increases the sketching rank conditions to
@@
Finally, we address the application of our framework to operations intrinsic to the tensor train format. An important implication of the property established by Theorem 3.10 is the quality of the TTStack embedding for its application to the Randomized TT-Rounding. The following theorem demonstrates that our sketch produces a rounding approximation that preserves the quasi-optimality of the randomized SVD.